# Hallucination Detection in LLMs Using Spectral Features of Attention Maps

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable performance across various tasks but remain prone to hallucinations. Detecting hallucinations is essential for safety-critical applications, and recent methods leverage attention map properties to this end, though their effectiveness remains limited. In this work, we investigate the spectral features of attention maps by interpreting them as adjacency matrices of graph structures. We propose the $\operatorname{LapEigvals}$ method, which utilises the top-$k$ eigenvalues of the Laplacian matrix derived from the attention maps as an input to hallucination detection probes. Empirical evaluations demonstrate that our approach achieves state-of-the-art hallucination detection performance among attention-based methods. Extensive ablation studies further highlight the robustness and generalisation of $\operatorname{LapEigvals}$, paving the way for future advancements in the hallucination detection domain.
@@
Jakub Binkowski1, Denis Janiak1, Albert Sawczyn1, Bogdan Gabrys2, Tomasz Kajdanowicz1 1Wroclaw University of Science and Technology, 2University of Technology Sydney, Correspondence: jakub.binkowski@pwr.edu.pl
@@ 1 Introduction
The recent surge of interest in Large Language Models (LLMs), driven by their impressive performance across various tasks, has led to significant advancements in their training, fine-tuning, and application to real-world problems. Despite progress, many challenges remain unresolved, particularly in safety-critical applications where the cost of errors is high. A significant issue is that LLMs are prone to hallucinations, i.e. generating "content that is nonsensical or unfaithful to the provided source content" (Farquhar et al., 2024; Huang et al., 2023). Since eliminating hallucinations is impossible (Lee, 2023; Xu et al., 2024), there is a pressing need for methods to detect when a model produces hallucinations. In addition, uncovering internal behaviour while studying hallucinations of LLMs might reveal significant progress in understanding their characteristics, fostering further development in the field. Recent studies have shown that hallucinations can be detected using internal states of the model, e.g., hidden states (Chen et al., 2024) or attention maps (Chuang et al., 2024a), and that LLMs can internally "know when they do not know" (Azaria and Mitchell, 2023; Orgad et al., 2025). We provide new insights showing that spectral features of attention maps coincide with hallucinations, and based on that observation, we introduce a novel method for detecting hallucinations.
As highlighted by (Barbero et al., 2024), attention maps can be viewed as weighted adjacency matrices of graphs. Building on this perspective, we performed statistical analysis and demonstrated that the eigenvalues of a Laplacian matrix derived from attention maps serve as good predictors of hallucinations. We propose the $\operatorname{LapEigvals}$ method, which utilises the top-$k$ eigenvalues of the Laplacian as input features of a probing model to detect hallucinations. We share full implementation in a public repository: https://github.com/graphml-lab-pwr/lapeig.
We summarise our contributions as follows:
@@
    Through extensive ablation studies, we demonstrate properties, robustness and generalisation of $\operatorname{LapEigvals}$ and suggest promising directions for further development.
@@ 2 Motivation
Figure 1: Visualisation of $p$-values from the two-sided Mann-Whitney U test for all layers and heads of Llama-3.1-8B across two feature types: $\operatorname{AttentionScore}$ and the $k=10$ Laplacian eigenvalues. These features were derived from attention maps collected when the LLM answered questions from the TriviaQA dataset. Higher $p$-values indicate no significant difference in feature values between hallucinated and non-hallucinated examples. For $\operatorname{AttnEigvals}$, $80\%$ of heads have $p<0.05$, while for Laplacian eigenvalues, this percentage is $91\%$. Therefore, Laplacian eigenvalues may be better predictors of hallucinations, as feature values across more heads exhibit statistically significant differences between hallucinated and non-hallucinated examples.
Considering the attention matrix as an adjacency matrix representing a set of Markov Chains, each corresponding to one layer of an LLM (Barbero et al., 2024) (Figure 2), we can leverage its spectral properties, as was done in many successful graph-based methods (Mohar, 1997; von Luxburg, 2007; Bruna et al., 2013; Topping et al., 2022). In particular, it was shown that graph Laplacian might help to describe several graph properties, like the presence of bottlenecks (Topping et al., 2022; Black et al., 2023). We hypothesise that hallucinations may be related to disturbance of information flow caused by some form of bottleneck.
To assess whether our hypothesis holds, we measured if graph spectral features provide a stronger coincidence with hallucinations than the previous attention-based method - $\operatorname{AttentionScore}$ (Sriramanan et al., 2024). We prompted an LLM with questions from the TriviaQA dataset (Joshi et al., 2017) and extracted attention maps, differentiating by layers and heads. We then computed the spectral features, i.e., the 10 largest eigenvalues of the Laplacian matrix from each head and layer. Further, we conducted a two-sided Mann-Whitney U test to compare whether Laplacian eigenvalues and the values of $\operatorname{AttentionScore}$ are different between hallucinated and non-hallucinated examples. Figure 1 shows $p$-values for all layers and heads, indicating that $\operatorname{AttentionScore}$ often results in higher $p$-values compared to Laplacian eigenvalues. Overall, we studied 6 datasets and 2 LLMs and found similar results, and present all results in Appendix A. Based on these findings, we propose leveraging top-$k$ Laplacian eigenvalues as features for a hallucination probe.
@@ 3 Method
In our method, we train a hallucination probe using only attention maps extracted during LLM inference, as illustrated in Figure 2. The attention map is a matrix containing attention scores for all tokens processed during inference, while the hallucination probe is a logistic regression model that uses features derived from attention maps as input. This work’s core contribution is using the top-$k$ eigenvalues of the Laplacian matrix as input features, which we detail below.
Denote $\mathbf{A}^{(l,h)}\in\mathbb{R}^{T\times T}$ as the attention map matrix for layer $l\in\{1\dotsc L\}$ and attention head $h\in\{1\dotsc H\}$, where $T$ is the total number of tokens generated by an LLM (including input tokens), $L$ the number of layers (transformer blocks), and $H$ the number of attention heads. The attention matrix is row-stochastic, meaning each row sums to 1 ($\sum_{j=0}^{T}\mathbf{A}^{(l,h)}_{:,j}=\mathbf{1}$). It is also lower triangular ($a^{(l,h)}_{ij}=1$ for all $j>i$) and non-negative ($a^{(l,h)}_{ij}\geq 0$ for all $i,j$). We can view $\mathbf{A}^{(l,h)}$ as a weighted adjacency matrix of a directed graph, where each node represents processed token, and each directed edge from token $i$ to token $j$ is weighted by the attention score, as depicted in Figure 2.
@@
$d^{(l,h)}_{ii}=\frac{\sum_{u}{a^{(l,h)}_{ui}}}{\sum_{v}{\mathbb{I}\left\{a^{(l,h)}_{vi}\neq 0\right\}}},$ (2)
where $\mathbb{I}$ is an indicator function, and $i,u,v\in\{1\dots T\}$ denote token indices. Intuitively, the resulting Laplacian for each processed token represents the average attention score to previous tokens reduced by the attention score to itself. Due to the structure of LLM inference, this formulation ensures that the Laplacian remains a lower triangular matrix. Hence, diagonal entries are eigenvalues, and there is no need for expensive eigendecomposition.
A recent study by (Orgad et al., 2025) has demonstrated that selecting the appropriate token and layer to take hidden states from is complex and significantly affects hallucination detection. Conversely, (Zhu et al., 2024) found that using the entire sequence of hidden states, rather than a single token, improves the detection performance. In addition, (Kim et al., 2024) showed that using information from all layers instead of one in isolation leads to better performance on this task. Motivated by these results, in our method, we consider all layers and all tokens.
Therefore, we take eigenvalues of $\mathbf{L}^{(l,h)}$, which are diagonal entries due to the lower triangularity of the matrix, and sort them:
@@
Finally, we take the top-$k$ largest values and concatenate them into a single feature vector $z$, where $k$ is a hyperparameter of our method:
@@
Since LLMs contain dozens of layers and heads, vector $z$ would suffer from large dimensionality. Thus, we decompose it to lower dimensionality using the PCA (Jolliffe and Cadima, 2016). We call our approach $\operatorname{LapEigvals}$.
@@ 4.1 Dataset construction
We use annotated QA datasets to construct the hallucination detection datasets and label incorrect LLM answers as hallucinations. To determine whether the generated answers were correct, we adopted the llm-as-judge approach (Zheng et al., 2023), as in previous studies (Orgad et al., 2025). Specifically, we prompted a large LLM to classify each response as either hallucination, non-hallucination, or rejected, where rejected indicates that it was unclear whether the answer was correct, e.g., the model refused to answer due to insufficient knowledge. Based on the manual qualitative inspection of several LLMs, we employed gpt-4o-mini (OpenAI et al., 2024) as the judge model since it provides the best trade-off between accuracy and cost.
For experiments, we selected 6 QA datasets previously utilised in the context of hallucination detection (Chen et al., 2024; Kossen et al., 2024; Chuang et al., 2024b; Mitra et al., 2024). Specifically, we used the validation set of NQOpen (Kwiatkowski et al., 2019), comprising $3610$ question-answer pairs, and the validation set of TriviaQA (Joshi et al., 2017), containing 7983 pairs. To evaluate our method on longer inputs, we employed the development set of CoQA (Reddy et al., 2019) and the rc.nocontext portion of the SQuADv2 (Rajpurkar et al., 2018) datasets, with 5928 and 9960 examples, respectively. Additionally, we incorporated the QA part of the HaluEval (Li et al., 2023) dataset, containing $10000$ examples, and the generation part of the TruthfulQA (Lin et al., 2022) benchmark with 817 examples. For TriviaQA, CoQA, and SQuADv2, we followed the same preprocessing procedure as (Chen et al., 2024).
We generate answers using 3 open-source LLMs: Llama-3.1-8B111hf.co/meta-llama/Llama-3.1-8B-Instruct and Llama-3.2-3B222hf.co/meta-llama/Llama-3.2-3B-Instruct (Grattafiori et al., 2024), Phi-3.5333hf.co/microsoft/Phi-3.5-mini-instruct (Abdin et al., 2024). We use two softmax temperatures for each LLM when decoding ($temp\in\{0.1,1.0\}$) and one prompt (showed on Listing 3). Overall, we evaluated hallucination detection probes on 6 LLM configurations and 6 QA datasets. We present the frequency of classes for answers from each configuration in Figure 9 (Appendix C).
@@ 4.2 Hallucination Probe
As a hallucination probe, we take a logistic regression model, using the implementation from scikit-learn (Pedregosa et al., 2011) with all parameters default, except for ${max\_iter=2000}$ and ${class\_weight={{}^{\prime\prime}balanced^{\prime\prime}}}$. For top-$k$ eigenvalues, we tested 5 values of $k\in\{5,10,20,50,100\}$ [^4] For datasets with examples having less than 100 tokens, we stop at $k=50$ and selected the result with the highest efficacy. All eigenvalues are projected with PCA onto 512 dimensions, except in per-layer experiments where there may be fewer than 512 features. In these cases, we apply PCA projection to match the input feature dimensionality, i.e., decorrelating them. As an evaluation metric, we use AUROC on the test split.
@@ 4.3 Baselines
Our method is a supervised approach to detect hallucinations solely from attention maps. For a fair comparison, we modify unsupervised $\operatorname{AttentionScore}$ (Sriramanan et al., 2024) to take log-determinants for each head as features instead of summing them. We also add original $\operatorname{AttentionScore}$ with the summation over heads for a reference. To evaluate the effectiveness of our proposed Laplacian eigenvalues, we also compare it to using raw attention maps and call it $\operatorname{AttnEigvals}$. Additionally, in Appendix D, we provide results for each approach but per-layer.
4.4 Implementation details
In our experiments, we used HuggingFace Transformers (Wolf et al., 2020), PyTorch (Ansel et al., 2024), and scikit-learn (Pedregosa et al., 2011). We utilised Pandas (team, 2020) and Seaborn (Waskom, 2021) for visualisations and analysis. To version data, we employed DVC (Kuprieiev et al., 2025). We acquired attention maps using a single Nvidia A40 with 40GB VRAM, remaining computations were carried out on CPU. To compute labels using the llm-as-judge approach, we leveraged OpenAI API. Detailed hyperparameter configurations and code to reproduce the experiments is available in the public Git repository.
@@ 5 Results
Table 1: Test AUROC for $\operatorname{LapEigvals}$ and several baseline methods. AUROC values were obtained in the single run of logistic regression training on features from a dataset generated with $temp=1.0$. We marked results for $\operatorname{AttentionScore}$ in gray as it is unsupervised approach, not directly comparable to the other ones. In bold, we highlight the best performance individually for each dataset and LLM. See Appendix D for extended results.
@@
CoQA HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B $\operatorname{AttentionScore}$ 0.493 0.589 0.556 0.538 0.532 0.541
Llama3.1-8B $\operatorname{AttnLogDet}$ 0.769 0.827 0.793 0.748 0.842 0.814
Llama3.1-8B $\operatorname{AttnEigvals}$ 0.782 0.819 0.790 0.768 0.843 0.833
Llama3.1-8B $\operatorname{LapEigvals}$ 0.830 0.874 0.827 0.791 0.889 0.829
Llama3.2-3B $\operatorname{AttentionScore}$ 0.509 0.588 0.546 0.530 0.515 0.581
Llama3.2-3B $\operatorname{AttnLogDet}$ 0.700 0.801 0.690 0.734 0.789 0.795
Llama3.2-3B $\operatorname{AttnEigvals}$ 0.724 0.819 0.694 0.749 0.804 0.723
Llama3.2-3B $\operatorname{LapEigvals}$ 0.812 0.828 0.693 0.757 0.832 0.787
Phi3.5 $\operatorname{AttentionScore}$ 0.520 0.541 0.594 0.504 0.540 0.554
Phi3.5 $\operatorname{AttnLogDet}$ 0.745 0.818 0.815 0.769 0.848 0.755
Phi3.5 $\operatorname{AttnEigvals}$ 0.771 0.829 0.798 0.782 0.850 0.802
Phi3.5 $\operatorname{LapEigvals}$ 0.821 0.836 0.826 0.795 0.872 0.777
Table 1 presents the results of our method compared to the baselines. $\operatorname{LapEigvals}$ achieved the best performance among all tested methods on 5 out of 6 datasets. Moreover, our method consistently performs well across all three LLM architectures. TruthfulQA was the only exception where $\operatorname{LapEigvals}$ was the second-best approach, yet it might stem from the small size of the dataset or severe class imbalance (depicted in Figure 9). In contrast, using eigenvalues of vanilla attention maps in $\operatorname{AttnEigvals}$ leads to worse performance, which suggests that transformation to Laplacian is the crucial step to uncover latent features of an LLM corresponding to hallucinations. In Appendix D, we show that $\operatorname{LapEigvals}$ consistently demonstrates a smaller generalisation gap, i.e., the difference between training and test performance is smaller for our method. While the $\operatorname{AttnLogDet}$ method performed poorly, it is fully unsupervised and should not be directly compared to other approaches. However, its supervised counterpart – $\operatorname{AttnLogDet}$ – remains inferior to methods based on spectral features, namely $\operatorname{LapEigvals}$ and $\operatorname{AttnEigvals}$. In Table 4 in Appendix 4, we present extended results, including per-layer and all-layers breakdowns, two temperatures used during answer generation, and a comparison between training and test AUROC.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable performance across various tasks but remain prone to hallucinations. Detecting hallucinations is essential for safety-critical applications, and recent methods leverage attention map properties to this end, though their effectiveness remains limited. In this work, we investigate the spectral features of attention maps by interpreting them as adjacency matrices of graph structures. We propose the $\operatorname{LapEigvals}$ method, which utilizes the top-$k$ eigenvalues of the Laplacian matrix derived from the attention maps as an input to hallucination detection probes. Empirical evaluations demonstrate that our approach achieves state-of-the-art hallucination detection performance among attention-based methods. Extensive ablation studies further highlight the robustness and generalization of $\operatorname{LapEigvals}$, paving the way for future advancements in the hallucination detection domain.
@@
Jakub Binkowski1, Denis Janiak1, Albert Sawczyn1 Bogdan Gabrys2, Tomasz Kajdanowicz1 1Wroclaw University of Science and Technology, 2University of Technology Sydney, Correspondence: jakub.binkowski@pwr.edu.pl
@@ 1 Introduction
The recent surge of interest in Large Language Models (LLMs), driven by their impressive performance across various tasks, has led to significant advancements in their training, fine-tuning, and application to real-world problems. Despite progress, many challenges remain unresolved, particularly in safety-critical applications with a high cost of errors. A significant issue is that LLMs are prone to hallucinations, i.e. generating "content that is nonsensical or unfaithful to the provided source content" (Farquhar et al., 2024; Huang et al., 2023). Since eliminating hallucinations is impossible (Lee, 2023; Xu et al., 2024), there is a pressing need for methods to detect when a model produces hallucinations. In addition, examining the internal behavior of LLMs in the context of hallucinations may yield important insights into their characteristics and support further advancements in the field. Recent studies have shown that hallucinations can be detected using internal states of the model, e.g., hidden states (Chen et al., 2024) or attention maps (Chuang et al., 2024a), and that LLMs can internally "know when they do not know" (Azaria and Mitchell, 2023; Orgad et al., 2025). We show that spectral features of attention maps coincide with hallucinations and, building on this observation, propose a novel method for their detection.
As highlighted by (Barbero et al., 2024), attention maps can be viewed as weighted adjacency matrices of graphs. Building on this perspective, we performed statistical analysis and demonstrated that the eigenvalues of a Laplacian matrix derived from attention maps serve as good predictors of hallucinations. We propose the $\operatorname{LapEigvals}$ method, which utilizes the top-$k$ eigenvalues of the Laplacian as input features of a probing model to detect hallucinations. We share full implementation in a public repository: https://github.com/graphml-lab-pwr/lapeigvals.
We summarize our contributions as follows:
@@
    Through extensive ablation studies, we demonstrate properties, robustness and generalization of $\operatorname{LapEigvals}$ and suggest promising directions for further development.
@@ 2 Motivation
Figure 1: Visualization of $p$-values from the two-sided Mann-Whitney U test for all layers and heads of Llama-3.1-8B across two feature types: $\operatorname{AttentionScore}$ and the k=10k{=}10 Laplacian eigenvalues. These features were derived from attention maps collected when the LLM answered questions from the TriviaQA dataset. Higher $p$-values indicate no significant difference in feature values between hallucinated and non-hallucinated examples. For $\operatorname{AttentionScore}$, $80\%$ of heads have $p<0.05$, while for Laplacian eigenvalues, this percentage is $91\%$. Therefore, Laplacian eigenvalues may be better predictors of hallucinations, as feature values across more heads exhibit statistically significant differences between hallucinated and non-hallucinated examples.
Considering the attention matrix as an adjacency matrix representing a set of Markov chains, each corresponding to one layer of an LLM (Wu et al., 2024) (see Figure 2), we can leverage its spectral properties, as was done in many successful graph-based methods (Mohar, 1997; von Luxburg, 2007; Bruna et al., 2013; Topping et al., 2022). In particular, it was shown that the graph Laplacian might help to describe several graph properties, like the presence of bottlenecks (Topping et al., 2022; Black et al., 2023). We hypothesize that hallucinations may arise from disruptions in information flow, such as bottlenecks, which could be detected through the graph Laplacian.
To assess whether our hypothesis holds, we computed graph spectral features and verified if they provide a stronger coincidence with hallucinations than the previous attention-based method - $\operatorname{AttentionScore}$ (Sriramanan et al., 2024). We prompted an LLM with questions from the TriviaQA dataset (Joshi et al., 2017) and extracted attention maps, differentiating by layers and heads. We then computed the spectral features, i.e., the 10 largest eigenvalues of the Laplacian matrix from each head and layer. Further, we conducted a two-sided Mann-Whitney U test (Mann and Whitney, 1947) to compare whether Laplacian eigenvalues and the values of $\operatorname{AttentionScore}$ are different between hallucinated and non-hallucinated examples. Figure 1 shows $p$-values for all layers and heads, indicating that $\operatorname{AttentionScore}$ often results in higher $p$-values compared to Laplacian eigenvalues. Overall, we studied 7 datasets and 5 LLMs and found similar results (see Appendix A). Based on these findings, we propose leveraging top-$k$ Laplacian eigenvalues as features for a hallucination probe.
@@ 3 Method
In our method, we train a hallucination probe using only attention maps, which we extracted during LLM inference, as illustrated in Figure 2. The attention map is a matrix containing attention scores for all tokens processed during inference, while the hallucination probe is a logistic regression model that uses features derived from attention maps as input. This work’s core contribution is using the top-$k$ eigenvalues of the Laplacian matrix as input features, which we detail below.
Denote $\mathbf{A}^{(l,h)}\in\mathbb{R}^{T\times T}$ as the attention map matrix for layer $l\in\{1\dotsc L\}$ and attention head $h\in\{1\dotsc H\}$, where $T$ is the total number of tokens generated by an LLM (including input tokens), $L$ the number of layers (transformer blocks), and $H$ the number of attention heads. The attention matrix is row-stochastic, meaning each row sums to 1 ($\sum_{j=0}^{T}\mathbf{A}^{(l,h)}_{:,j}=\mathbf{1}$). It is also lower triangular ($a^{(l,h)}_{ij}=0$ for all $j>i$) and non-negative ($a^{(l,h)}_{ij}\geq 0$ for all $i,j$). We can view $\mathbf{A}^{(l,h)}$ as a weighted adjacency matrix of a directed graph, where each node represents processed token, and each directed edge from token $i$ to token $j$ is weighted by the attention score, as depicted in Figure 2.
@@
$d^{(l,h)}_{ii}=\frac{\sum_{u}{a^{(l,h)}_{ui}}}{T-i},$ (2)
where $i,u\in\{0,\dots,(T-1)\}$ denote token indices. The Laplacian defined this way is bounded, i.e., $\mathbf{L}^{(l,h)}_{ij}\in\left[-1,1\right]$ (see Appendix B for proofs). Intuitively, the resulting Laplacian for each processed token represents the average attention score to previous tokens reduced by the attention score to itself. As eigenvalues of the Laplacian can summarize information flow in a graph (von Luxburg, 2007; Topping et al., 2022), we take eigenvalues of $\mathbf{L}^{(l,h)}$, which are diagonal entries due to the lower triangularity of the Laplacian matrix, and sort them:
@@
Recently, (Zhu et al., 2024) found features from the entire token sequence, rather than a single token, improving hallucination detection. Similarly, (Kim et al., 2024) demonstrated that information from all layers, instead of a single one in isolation, yields better results on this task. Motivated by these findings, our method uses features from all tokens and all layers as input to the probe. Therefore, we take the top-$k$ largest values from each head and layer and concatenate them into a single feature vector $z$, where $k$ is a hyperparameter of our method:
@@
Since LLMs contain dozens of layers and heads, the probe input vector $z\in\mathbb{R}^{L\cdot H\cdot k}$ can still be high-dimensional. Thus, we project it to a lower dimensionality using PCA (Jolliffe and Cadima, 2016). We call our approach $\operatorname{LapEigvals}$.
@@ 4 Experimental setup
The overview of the methodology used in this work is presented in Figure 3. Next, we describe each step of the pipeline in detail.
@@ 4.1 Dataset construction
We use annotated QA datasets to construct the hallucination detection datasets and label incorrect LLM answers as hallucinations. To assess the correctness of generated answers, we followed prior work (Orgad et al., 2025) and adopted the llm-as-judge approach (Zheng et al., 2023), with the exception of one dataset where exact match evaluation against ground-truth answers was possible. For llm-as-judge, we prompted a large LLM to classify each response as either hallucination, non-hallucination, or rejected, where rejected indicates that it was unclear whether the answer was correct, e.g., the model refused to answer due to insufficient knowledge. Based on the manual qualitative inspection of several LLMs, we employed gpt-4o-mini (OpenAI et al., 2024) as the judge model since it provides the best trade-off between accuracy and cost. To confirm the reliability of the labels, we additionally verified agreement with the larger model, gpt-4.1, on Llama-3.1-8B and found that the agreement between models falls within the acceptable range widely adopted in the literature (see Appendix F).
For experiments, we selected 7 QA datasets previously utilized in the context of hallucination detection (Chen et al., 2024; Kossen et al., 2024; Chuang et al., 2024b; Mitra et al., 2024). Specifically, we used the validation set of NQ-Open (Kwiatkowski et al., 2019), comprising 3,6103{,}610 question-answer pairs, and the validation set of TriviaQA (Joshi et al., 2017), containing 7,9837{,}983 pairs. To evaluate our method on longer inputs, we employed the development set of CoQA (Reddy et al., 2019) and the rc.nocontext portion of the SQuADv2 (Rajpurkar et al., 2018) datasets, with 5,9285{,}928 and 9,9609{,}960 examples, respectively. Additionally, we incorporated the QA part of the HaluEvalQA (Li et al., 2023) dataset, containing 10,00010{,}000 examples, and the generation part of the TruthfulQA (Lin et al., 2022) benchmark with $817$ examples. Finally, we used test split of GSM8k dataset Cobbe et al. (2021) with 1,3191{,}319 grade school math problems, evaluated by exact match against labels. For TriviaQA, CoQA, and SQuADv2, we followed the same preprocessing procedure as (Chen et al., 2024).
We generate answers using 5 open-source LLMs: Llama-3.1-8B111hf.co/meta-llama/Llama-3.1-8B-Instruct and Llama-3.2-3B222hf.co/meta-llama/Llama-3.2-3B-Instruct (Grattafiori et al., 2024), Phi-3.5333hf.co/microsoft/Phi-3.5-mini-instruct (Abdin et al., 2024), Mistral-Nemo444hf.co/mistralai/Mistral-Nemo-Instruct-2407 (Mistral AI Team and NVIDIA, 2024), Mistral-Small-24B555hf.co/mistralai/Mistral-Small-24B-Instruct-2501 (Mistral AI Team, 2025). We use two softmax temperatures for each LLM when decoding ($temp\in\{0.1,1.0\}$) and one prompt (for all datasets we used a prompt in Listing 3, except for GSM8K in Listing 5). Overall, we evaluated hallucination detection probes on 10 LLM configurations and 7 QA datasets. We present the frequency of classes for answers from each configuration in Figure 9 (Appendix E).
@@ 4.2 Hallucination Probe
As a hallucination probe, we take a logistic regression model, using the implementation from scikit-learn (Pedregosa et al., 2011) with all parameters default, except for max_iter=2000{max\_iter{=}2000} and $class\_weight{=}{{}^{\prime\prime}balanced^{\prime\prime}}$}. For top-$k$ eigenvalues, we tested 5 values of $k\in\{5,10,20,50,100\}$ [^6] For datasets with examples having less than 100 tokens, we stop at k=50k{=}50 and selected the result with the highest efficacy. All eigenvalues are projected with PCA onto 512 dimensions, except in per-layer experiments where there may be fewer than 512 features. In these cases, we apply PCA projection to match the input feature dimensionality, i.e., decorrelating them. As an evaluation metric, we use AUROC on the test split (additional results presenting Precision and Recall are reported in Appendix G.1).
@@ 4.3 Baselines
Our method is a supervised approach for detecting hallucinations using only attention maps. For a fair comparison, we adapt the unsupervised $\operatorname{AttentionScore}$ (Sriramanan et al., 2024) by using log-determinants of each head’s attention scores as features instead of summing them, and we also include the original $\operatorname{AttentionScore}$, computed as the sum of log-determinants over heads, for reference. To evaluate the effectiveness of our proposed Laplacian eigenvalues, we compare them to the eigenvalues of raw attention maps, denoted as $\operatorname{AttnEigvals}$. Extended results for each approach on a per-layer basis are provided in Appendix G.2, while Appendix G.4 presents a comparison with a method based on hidden states. Implementation and hardware details are provided in Appendix C.
@@ 5 Results
Table 1: Test AUROC for $\operatorname{LapEigvals}$ and several baseline methods. AUROC values were obtained in a single run of logistic regression training on features from a dataset generated with temp=1.0temp{=}1.0. We mark results for $\operatorname{AttentionScore}$ in gray as it is an unsupervised approach, not directly comparable to the others. In bold, we highlight the best performance individually for each dataset and LLM. See Appendix G for extended results.
@@
CoQA GSM8K HaluevalQA NQOpen SQuADv2 TriviaQA TruthfulQA
Llama3.1-8B $\operatorname{AttentionScore}$ 0.493 0.720 0.589 0.556 0.538 0.532 0.541
Llama3.1-8B $\operatorname{AttnLogDet}$ 0.769 0.826 0.827 0.793 0.748 0.842 0.814
Llama3.1-8B $\operatorname{AttnEigvals}$ 0.782 0.838 0.819 0.790 0.768 0.843 0.833
Llama3.1-8B $\operatorname{LapEigvals}$ 0.830 0.872 0.874 0.827 0.791 0.889 0.829
Llama3.2-3B $\operatorname{AttentionScore}$ 0.509 0.717 0.588 0.546 0.530 0.515 0.581
Llama3.2-3B $\operatorname{AttnLogDet}$ 0.700 0.851 0.801 0.690 0.734 0.789 0.795
Llama3.2-3B $\operatorname{AttnEigvals}$ 0.724 0.768 0.819 0.694 0.749 0.804 0.723
Llama3.2-3B $\operatorname{LapEigvals}$ 0.812 0.870 0.828 0.693 0.757 0.832 0.787
Phi3.5 $\operatorname{AttentionScore}$ 0.520 0.666 0.541 0.594 0.504 0.540 0.554
Phi3.5 $\operatorname{AttnLogDet}$ 0.745 0.842 0.818 0.815 0.769 0.848 0.755
Phi3.5 $\operatorname{AttnEigvals}$ 0.771 0.794 0.829 0.798 0.782 0.850 0.802
Phi3.5 $\operatorname{LapEigvals}$ 0.821 0.885 0.836 0.826 0.795 0.872 0.777
Mistral-Nemo $\operatorname{AttentionScore}$ 0.493 0.630 0.531 0.529 0.510 0.532 0.494
Mistral-Nemo $\operatorname{AttnLogDet}$ 0.728 0.856 0.798 0.769 0.772 0.812 0.852
Mistral-Nemo $\operatorname{AttnEigvals}$ 0.778 0.842 0.781 0.761 0.758 0.821 0.802
Mistral-Nemo $\operatorname{LapEigvals}$ 0.835 0.890 0.833 0.795 0.812 0.865 0.828
Mistral-Small-24B $\operatorname{AttentionScore}$ 0.516 0.576 0.504 0.462 0.455 0.463 0.451
Mistral-Small-24B $\operatorname{AttnLogDet}$ 0.766 0.853 0.842 0.747 0.753 0.833 0.735
Mistral-Small-24B $\operatorname{AttnEigvals}$ 0.805 0.856 0.848 0.751 0.760 0.844 0.765
Mistral-Small-24B $\operatorname{LapEigvals}$ 0.861 0.925 0.882 0.791 0.820 0.876 0.748
Table 1 presents the results of our method compared to the baselines. $\operatorname{LapEigvals}$ achieved the best performance among all tested methods on 6 out of 7 datasets. Moreover, our method consistently performs well across all 5 LLM architectures ranging from 3 up to 24 billion parameters. TruthfulQA was the only exception where $\operatorname{LapEigvals}$ was the second-best approach, yet it might stem from the small size of the dataset or severe class imbalance (depicted in Figure 9). In contrast, using eigenvalues of vanilla attention maps in $\operatorname{AttnEigvals}$ leads to worse performance, which suggests that transformation to Laplacian is the crucial step to uncover latent features of an LLM corresponding to hallucinations. In Appendix G, we show that $\operatorname{LapEigvals}$ consistently demonstrates a smaller generalization gap, i.e., the difference between training and test performance is smaller for our method. While the $\operatorname{AttentionScore}$ method performed poorly, it is fully unsupervised and should not be directly compared to other approaches. However, its supervised counterpart – $\operatorname{AttnLogDet}$ – remains inferior to methods based on spectral features, namely $\operatorname{AttnEigvals}$ and $\operatorname{LapEigvals}$. In Table 6 in Appendix G.2, we present extended results, including per-layer and all-layers breakdowns, two temperatures used during answers generation, and a comparison between training and test AUROC. Moreover, compared to probes based on hidden states, our method performs best in most of the tested settings, as shown in Appendix G.4.